OpenStudy (loser66):

If u =(x1,x2) , v = (y1,y2) f(u,v) = 2x1y2-3x2y1 how to show f is a bilinear form? Please help

OpenStudy (ybarrap):

Hi @loser! Did you check each of the sum tests? http://en.wikipedia.org/wiki/Bilinear_form

OpenStudy (loser66):

I don't know what 's wrong with me. My brain is on and off with the stuff. Sometime, it's quite trivial to me but just few hours later, it disappears as if I am a dummy one. :(

OpenStudy (loser66):

@dan815 can you help me?

OpenStudy (loser66):

@wio @dumbcow

OpenStudy (anonymous):

Prove the properties on the wikipedia page.

OpenStudy (loser66):

I don't know how. :(

OpenStudy (anonymous):

Show that \[ f(u+w,v) = f(u,v) + f(w,v) \]first

OpenStudy (anonymous):

let \(w=\langle z_1,z_2\rangle\)

OpenStudy (anonymous):

If you are tired, get some rest. This is too trivial for me to do it for you.

OpenStudy (loser66):

Please!! I spend whole day to understand the stuff. I know how to prove in symbolic form only. :(

OpenStudy (loser66):

I mean I can put 2 cases in 1like f(u1,u2)(v1,v2) = f(u1, v1) + f(u1,v2)+ f(u2,v1)+ f(u2,v2) but ...blank then

OpenStudy (anonymous):

n \[ u+w = \langle x_1+z_1,x_2+z_2\rangle \]

OpenStudy (anonymous):

That's my limit.

OpenStudy (loser66):

Thanks anyway!!:( let me try later. So tired now

OpenStudy (ikram002p):

w=(z1,z2) If u =(x1,x2) , v = (y1,y2) f(u,v) = 2x1y2-3x2y1 f(u+v,w)=f(u,w)+f(v,w) check if 2(x1+y1)(z2)-3(x2+y2)(z1)=(2x1z2-3x2z1)+(2y1z2-3y2z1) 2x1z2+2y1z2-3x2z1-3y2z1=(2x1z2-3x2z1)+(2y1z2-3y2z1) (2x1z2-3x2z1)+(2y1z1-3y2z1)=(2x1z2-3x2z1)+(2y1z2-3y2z1) so f(u+v,w)=f(u,w)+f(v,w) is true

OpenStudy (ikram002p):

f(u,v) = 2x1y2-3x2y1 next one f(u ,v+w)=f(u,v)+f(u,w) the same thing check if 2x1(y2+z2)-3x2(y1+z1)=(2x1y2-3x2y1)+(2x1z2-3x2z1) f(u ,v+w)=2x1(y2+z2)-3x2(y1+z1) =2x1y2+2x1z2-3x2y1-3x2z1 =(2x1y2-3x2y1)+(2x1z2-3x2z1) =f(u,v)+f(u,w) so f(u ,v+w)=f(u,v)+f(u,w) is true

OpenStudy (ikram002p):

the last one f(u,v) = 2x1y2-3x2y1 f(λu, v) = f(u, λv) = λf(u, v) case 1 check f(λu, v) = f(u, λv) 2(λx1)y2-3(λ2)y1=2x1(λy2)-3x2(λy1) f(λu, v) = 2(λx1)y2-3(λ2)y1 =2 λx1 y2-3 λx2 y1 =2x1 λy2 -3x2 λy1 =2x1(λy2)-3x2(λy1) = f(u, λv) case 2 f(λu, v) = λf(u, v) 2(λx1)y2-3(λ2)y1=λ(2x1 y2 -3x2 y1) f(λu, v) = 2(λx1)y2-3(λ2)y1 =2 λx1 y2-3 λx2 y1 =λ(2x1 y2 -3x2 y1) =λf(u, v) so f(λu, v) = f(u, λv) = λf(u, v) is true

OpenStudy (ybarrap):

Wow @ikram002p ! You are awesome!!

OpenStudy (ikram002p):

^^ np !